OpenStudy (anonymous):

What are the x- and y-intercepts for the graph of 5x – 2y = 20?

OpenStudy (anonymous):

To find the x intercept, sub in 0 for y 5x - 2y = 20 5x - 2(0) = 20 5x = 20 x = 4 so your x intercept is 4, or it could be written like this (4,0) To find the y intercept, sub in 0 for x 5x - 2y = 20 5(0) - 2y = 20 -2y = 20 y = -10 so your y intercept is -10, or (0,-10)

OpenStudy (anonymous):

thank you so much